OpenStudy (kyhannah):

How do I find the minimum value of -3+4cos[5pi/6(x+4)]

OpenStudy (dumbcow):

min = vertical shift - amplitude min = -3 -4 = -7

OpenStudy (kyhannah):

o: sweet thanks! :D i'll keep that formula in mind <3

OpenStudy (dumbcow):

yw, you can get same result too by noticing that cos has min value of -1 ---> -3 + 4(-1) = -7

OpenStudy (kyhannah):

ah ok :)